# Associative Transformer

### v3

@@ Abstract
Emerging from the pairwise attention in conventional Transformers, there is a growing interest in sparse attention mechanisms that align more closely with localized, contextual learning in the biological brain. Existing studies such as the Coordination method employ iterative cross-attention mechanisms with a bottleneck to enable the sparse association of inputs. However, these methods are parameter inefficient and fail in more complex relational reasoning tasks. To this end, we propose Associative Transformer (AiT) to enhance the association among sparsely attended input patches, improving parameter efficiency and performance in relational reasoning tasks. AiT leverages a learnable explicit memory, comprised of various specialized priors, with a bottleneck attention to facilitate the extraction of diverse localized features. Moreover, we propose a novel associative memory-enabled patch reconstruction with a Hopfield energy function. The extensive experiments in four image classification tasks with three different sizes of AiT demonstrate that AiT requires significantly fewer parameters and attention layers while outperforming Vision Transformers and a broad range of sparse Transformers. Additionally, AiT establishes new SOTA performance in the Sort-of-CLEVR dataset, outperforming the previous Coordination method.
Machine Learning, ICML
@@ 1 Introduction
Transformer models use pairwise attention to establish correlations among disparate segments of input information (Vaswani et al., 2017; Dosovitskiy et al., 2020). Emerging from the pair-wise attention, there is a growing interest in leveraging sparse interactions that align more closely with the localized, contextual learning in the biological brain. This sparsity attribute has demonstrated advantages in enhancing learning performance and efficiency (Brooks, 1991; Greff et al., 2020). Sparse knowledge association can find resonance with the neuroscientific grounding of the Global Workspace Theory (GWT) (Baars, 1988; Dehaene et al., 1998; VanRullen & Kanai, 2020; Juliani et al., 2022). GWT explains a fundamental cognitive architecture for working memory in the brain where diverse specialized modules compete to write information into a shared workspace through a communication bottleneck. The bottleneck facilitates the processing of content-addressable information using attention guided by contents in the shared workspace (Awh et al., 2006; Gazzaley & Nobre, 2012).
@@
A bottleneck guides models to generalize in a manner consistent with the underlying data distribution through inductive biases of sparsity (Baxter, 2000; Goyal & Bengio, 2022), resulting in superior performance in tasks such as relational reasoning. In this regard, the Coordination method (Goyal et al., 2022) is an initial attempt to assess the effectiveness of such a bottleneck in Transformers. Unfortunately, its design is parameter inefficient and often fails in more complex relational reasoning tasks. In this work, we aim to enhance sparse attention capability by leveraging low-rank explicit memory to learn a diverse set of priors to guide the attention. Drawing inspiration from associative memory operating on Hebbian learning in the hippocampus, we utilize continuous Hopfield networks (Hopfield, 2007; Ramsauer et al., 2021) to reconstruct patches from the explicit memory. The associative memory enhances the effective association of sparsely attended historical inputs.
@@
To this end, we propose Associative Transformer (AiT), a sparse representation learner augmented by associative memory. AiT learns explicit memory for each attention layer to facilitate the extraction of localized features at different abstraction levels. The explicit memory naturally evolves into specialized priors through bottleneck attention, with each focusing on a specific spatial relation for a task. We further propose a bottleneck attention balance loss to encourage diversity among the memory elements. Furthermore, the acquired priors could function as distinct attractors within the associative memory of Hopfield networks, facilitating input reconstruction from memory. Extensive empirical results demonstrate that the emerging specialization of priors and the reconstruction within associative memory could significantly enhance parameter efficiency and model performance across a broad spectrum of tasks.
Overall, the main contributions of this work are as follows:
  1. 1.
@@
    We propose a novel associative memory-enhanced Transformer called the Associative Transformer (AiT). AiT tackles inefficient sparse attention in conventional Transformers by leveraging naturally emerging specialized priors for guided attention and patch reconstruction within associative memory (Section 3.3).
  2. 2.
@@
    The learned priors function as attractors of Hopfield networks, facilitating information correlation and retrieval from historical inputs. To the best of our knowledge, this is the first study of leveraging Hopfield networks in the sparse attention mechanism of Transformers (Section 3.4).
  3. 3.
@@
    Extensive experiments including an ablation study demonstrate enhanced performance and parameter efficiency in various image classification and relational reasoning tasks. We show that even the smallest AiT-Small model surpasses the SOTA Coordination method in the relational reasoning tasks (Section 4.2, 4.7).
@@ 2 Related Work
This section provides a summary of the most relevant recent research on sparse attention, latent memory-enabled Transformers, and external memory mechanisms. We further investigate the relatedness of these studies to the different properties of the Global Workspace Theory in Appendix B.
@@
Studies on sparse Transformers have explored consolidating latent memory to extract localized, contextual representations from inputs (Lee et al., 2019; Gupta & Berant, 2020; Ma et al., 2021; Jaegle et al., 2021; Goyal et al., 2022; Jaegle et al., 2022). For example, Perceiver (Jaegle et al., 2021, 2022) utilized iterative cross-attention with a latent array and latent transformation to capture dependencies across input data. Set Transformer (Lee et al., 2019) and Linear Unified Nested Attention (Ma et al., 2021) also employed iterative cross-attention but without using latent transformation. Other methods, like Blockwise Self-Attention, relied on a strong assumption of predefined modularization for attention specialization (Qiu et al., 2020). In contrast, our method learns such modularization and attention specialization through end-to-end training. Unlike the vast majority of methods that employ latent memory, our proposed method leverages both explicit memory and associative memory for increased capacity and generality.
In a similar approach to constructing sparse Transformers based on GWT, the Coordination method (Goyal et al., 2022) introduced a bottleneck into cross-attention mechanisms to encourage neural module specialization. The results indicated that the competition via the bottleneck could contribute to a small set of specialized priors in latent memory to facilitate relational learning. However, the number of priors was limited to fewer than 10, all with the same dimension as the input representation. The evaluation was also restricted to simple tasks like the Triangle dataset. On the contrary, our method employs low-rank explicit memory to learn a significantly larger and richer set of priors (up to 128 priors learned from a pool of 32.8k patches). Importantly, the Coordination method relies on iterative cross-attentions, while our work focuses on a novel learning method of associative memory-enabled sparse attention.
This work is also related to modular neural networks and mixture of experts (Zuo et al., 2022; Allingham et al., 2022) in terms of competition in the shared workspace. Separating the information processing within Transformers into distinct components, depending on the input, has shown advantages in adding more flexibility in data processing.
@@
External memory such as tape storage and associative memory, has been employed in deep neural networks (Graves et al., 2014; Krotov & Hopfield, 2016; Gülçehre et al., 2018), with recent studies exploring the potential use of Hopfield networks and their modern variants (Hopfield, 2007; Demircigil et al., 2017; Ramsauer et al., 2021). In this work, we incorporate Hopfield networks as an integral element in the sparse attention mechanism of the proposed Associative Transformer. This approach is fundamentally different from the previous studies such as Energy Transformer (Hoover et al., 2023) that mainly focused on utilizing Hopfield networks independently of the attention mechanism.
@@ 3 Associative Transformer
Figure 1: The scheme of the Associative Transformer (AiT). (a) In a global workspace layer, the input $\mathbb{R}^{B\times N\times E}$ is squashed into vectors $\mathbb{R}^{(B\times N)\times E}$. The squashed representations are projected to a latent space of dimension $D<<E$ and are sparsely selected to update the explicit memory via a fixed bottleneck $k<<(B\times N)$. The Hopfield network utilizes the memory to reconstruct the input, where a learnable linear transformation (LT) scales the memory contents back to the input dimension $E$. (b) The Associative Transformer block consists of self attention, feed-forward layers, and the global workspace layer. Compared to Vision Transformer (ViT), leveraging the global workspace layer enhances the layer efficiency of the model. A shallower 6-layer AiT is shown to outperform a 12-layer ViT in Table 4.2.
This section discusses the essential building blocks of Associative Transformer (AiT), where we devise an associative memory-enabled attention layer called the Global Workspace Layer (GWL). The GWL comprises the low-rank explicit memory, the bottleneck attention, and the associative memory based on Hopfield networks (Figure 1).
@@ 3.1 Vision Transformers for Classification Tasks
Vision Transformers (ViT) tackle image classification tasks by processing sequences of image patches. The pre-processing layer partitions an image into non-overlapping patches. Let $x\in\mathbb{R}^{H\times W\times C}$ be an input from $X$, where $(H,W)$ is the resolution of the image and $C$ is the number of channels. $x$ is separated into a sequence of patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(P,P)$ is the resolution of each image patch and $N=\frac{HW}{P^{2}}$ is the number of patches. These patches are mapped to embeddings $v_{p}\in\mathbb{R}^{N\times E}$ with the linear projection. ViT leverages self-attention to map a query and a set of key-value pairs to an output. The patch embeddings are used to obtain the query, key, and value based on linear transformations $W^{Q}\in\mathbb{R}^{E\times D},\,W^{K}\in\mathbb{R}^{E\times D},\text{ and, }W^{V}\in\mathbb{R}^{E\times D}$. The output of an individual attention head is a weighted sum of the values, $\mbox{softmax}(\frac{W^{Q}_{i}v(W^{K}_{i}v)^{T}}{\sqrt{D}})\,W^{V}_{i}v$.
@@ 3.2 Low-Rank Priors in Explicit Memory
The Global Workspace Layer enables an efficient writing and reading mechanism in explicit memory. The writing operation allows the explicit memory to learn a set of low-rank priors from different input patches through cross-attention. In particular, we employ a squash layer to concatenate patch representations within the entire batch $V\in\mathbb{R}^{B\times N\times E}$ into vectors $V\in\mathbb{R}^{(B\times N)\times E}$, fostering association among patches across samples. In practice, the number of patches $B\times N$ varies depending on the batch size $B$. However, a bottleneck with a fixed capacity $k$ could restrict the number of attended patches to a constant value (Section 3.3).
The output of the squash layer is attended to by the multi-head cross-attention using the contents of the explicit memory as queries. The explicit memory aims to learn $M$ low-rank priors $\gamma=\mathbb{R}^{M\times D}$ where $D$ is the dimension of the prior, $D<<E$. The lower dimension is obtained through a downscaling linear transformation. We found that a rank of 32 and initialization with a Gaussian distribution works well in most cases. Priors are general assumptions about patches, such as the aggregated representations from different patches of the same attribute. The priors are used as various queries to compute cross-attentions that extract different sets of patches and update the explicit memory.
We devise a tailored cross-attention mechanism to update the explicit memory using the squash layer’s output $\Xi^{t}=V^{t}\in\mathbb{R}^{(B\times N)\times E}$. Notably, in the cross-attention, the query is a function of the current memory content $\gamma^{t}=\{\gamma^{t}_{i}\}_{i=1}^{M}$. The key and value are functions of the squashed input $\Xi^{t}$. The attention scores for head $i$ are computed by $\text{A}^{t}_{i}(\gamma^{t},\Xi^{t})=\text{softmax}(\frac{\gamma^{t}W^{Q}_{i,t}(\Xi^{t}W^{K}_{i,t})^{T}}{\sqrt{D}})$. This is the case of soft attention with limited constraints on the bottleneck capacity.
@@ 3.3 Bottleneck Attention with a Limited Capacity
We discuss the case of limiting the cross-attention capacity via top-$k$ hard attention. The hard attention constrains the number of patches that can be attended to at any given time, fostering competition among patches and ensuring the selection of the most relevant patches. Notably, we select the patches with the top-$k$ highest attention scores from $A^{t}_{i}$, $\text{head}^{t}_{i}=\text{top-}k(\text{A}^{t}_{i})\Xi^{t}W^{V}_{t}$. Additionally, to ensure a stable update of the explicit memory across different time steps, we employ layer normalization and the Exponentially Weighted Moving Average (EWMA) method as follows:
$\hat{\gamma}^{t}=\text{LN}(\text{Concat}(\text{head}^{t}_{1},\dots,\text{head}^{t}_{A})W^{O}),$ (2)
$\gamma^{t+1}=(1-\alpha)\cdot\gamma^{t}+\alpha\cdot\hat{\gamma}^{t},$ (3)
$\gamma^{t+1}=\frac{{\gamma^{t+1}}}{{\sqrt{\sum_{{j=1}}^{M}(\gamma^{t+1}_{j})^{2}}}},$
where LN is the layer normalization and $\alpha$ is a smoothing factor determining the decay rate of older observations, which is usually a small value such as 0.1. EWMA ensures stable memory updates by accumulating both old $\gamma^{t}$ and new memories $\hat{\gamma}^{t}$. During evaluation, we found that using different batch sizes has little influence on the test performance. This can be attributed to the small decay rate of the memory update.
@@
The sparsity induced by the bottleneck attention contributes to the emergence of specialized priors. However, cascading multiple bottleneck attentions could lead to difficulty in efficiently forming specialized priors (Figure 7). As information flows through multiple global workspace layers, the representations become diluted, necessitating a mechanism to counteract this inherent loss of input specificity. To overcome this challenge, we propose the bottleneck attention balance loss to encourage a more diverse selection of patches by the bottleneck attention. The bottleneck attention balance loss $\ell_{\text{bottleneck}}$ comprises two components, i.e., the accumulative attention score $\ell_{\text{{importance}}}$ and the chosen instances $\ell_{\text{{loads}}}$. The loss is computed for each patch position $l\in\{1,2,\dots,B\times N\}$ in the squash layer output. We derive the normalized variances across positions as follows:
$\ell_{\text{importance}_{i,l}}=\sum_{j=1}^{M}\text{A}^{t}_{i,j,l},$ (4)
$\ell_{\text{loads}_{i,l}}=\sum_{j=1}^{M}(\text{A}^{t}_{i,j,l}>0),$ (5)
$\ell_{\text{bottleneck}_{i}}=\frac{{\text{Var}(\{\ell_{\text{importance}_{i,l}}}\}_{l=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{l=1}^{B\times N}\ell_{\text{importance}_{i,l}}})^{2}+\epsilon}+$ (6)
$\frac{{\text{Var}(\{\ell_{\text{loads}_{i,l}}}\}_{l=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{l=1}^{B\times N}\ell_{\text{loads}_{i,l}}})^{2}+\epsilon},$
where $\text{A}^{t}_{i,j,l}$ is the attention score of the position $l$ for the $j$th memory slot of head $i$, Var($\cdot$) denotes the variance, and $\epsilon$ is a small value to avoid division by zero.
The losses for the different heads are summed up and added to the classification loss, $\ell=\ell_{\text{class}}+\sigma\cdot\sum_{i=1}^{A}\ell_{\text{bottleneck}_{i}},$ where $\sigma$ is a coefficient to adjust the ratio between the classification loss and the bottleneck attention balance loss.

### v4

@@ Abstract
Emerging from the pairwise attention in conventional Transformers, there is a growing interest in sparse attention mechanisms that align more closely with localized, contextual learning in the biological brain. Existing studies such as the Coordination method employ iterative cross-attention mechanisms with a bottleneck to enable the sparse association of inputs. However, these methods are parameter inefficient and fail in more complex relational reasoning tasks. To this end, we propose Associative Transformer (AiT) to enhance the association among sparsely attended input tokens, improving parameter efficiency and performance in various vision tasks such as classification and relational reasoning. AiT leverages a learnable explicit memory comprising specialized priors that guide bottleneck attentions to facilitate the extraction of diverse localized tokens. Moreover, AiT employs an associative memory-based token reconstruction using a Hopfield energy function. The extensive empirical experiments demonstrate that AiT requires significantly fewer parameters and attention layers outperforming a broad range of sparse Transformer models. Additionally, AiT outperforms the SOTA sparse Transformer models including the Coordination method on the Sort-of-CLEVR dataset.
@@ 1 Introduction
Transformer models use pairwise attention to establish correlations among disparate segments of input information [1, 2]. Emerging from the pair-wise attention, there is a growing interest in leveraging sparse interactions for localized, contextual learning. This sparsity attribute has demonstrated advantages in enhancing learning performance and efficiency [3, 4]. Sparse knowledge association can find resonance with the neuroscientific grounding of the Global Workspace Theory (GWT) [5, 6, 7, 8]. GWT explains a fundamental cognitive architecture for working memory where diverse specialized modules compete to write information into a shared workspace through a communication bottleneck. The bottleneck facilitates the processing of content-addressable information using attention guided by contents in the shared workspace [9, 10].
@@
A bottleneck guides models to generalize in a manner consistent with the underlying data distribution through inductive biases of sparsity [11, 12], resulting in superior performance. In this regard, the Coordination method [13] is an initial attempt to assess the effectiveness of such a bottleneck in Transformers. Unfortunately, its design is parameter inefficient and often fails in more complex relational reasoning tasks [14, 15, 16, 17, 18, 19, 20]. In this work, we aim to enhance sparse attention capability of Transformers by leveraging low-rank explicit memory to learn a diverse set of priors to guide the attention. Furthermore, drawing inspiration from associative memory based on Hebbian learning, we utilize continuous Hopfield networks [21, 22] to reconstruct input tokens from the explicit memory. The associative memory enhances the effective association of sparsely attended historical inputs.
@@
To this end, we propose Associative Transformer (AiT), a sparse representation learner. AiT learns explicit memory for each attention layer to facilitate the extraction of localized features at different abstraction levels. The explicit memory evolves into specialized priors through the bottleneck attention, with each focusing on a specific spatial relation of input tokens. We further propose a bottleneck attention balance loss to encourage diversity among the learned priors. Furthermore, the acquired priors function as distinct attractors within the associative memory of Hopfield networks, facilitating token reconstruction from memory. Extensive empirical results demonstrate that the emerging specialization of priors and the reconstruction within associative memory in AiT could significantly enhance parameter efficiency and model performance across a broad spectrum of tasks.
Overall, our main contributions are three-fold:
@@
1) We propose the Associative Transformer (AiT) to tackle inefficient sparse attention mechanisms in conventional Transformers, leveraging emerging specialized priors for guided bottleneck attention and token reconstruction within associative memory (Section 3.3).
@@
2) The learned priors function as attractors of Hopfield networks, facilitating information correlation and retrieval from historical inputs. To the best of our knowledge, this is the first study on consolidating Hopfield networks with the sparse attention mechanism (Section 3.4).
@@
3) Extensive experiments including an ablation study demonstrate enhanced parameter efficiency of AiT in various classification and relational reasoning tasks, outperforming the Coordination method (Section 4).
@@ 2 Related work
This section provides a summary of the most relevant recent work on sparse attention, latent memory-enabled Transformers, and external memory mechanisms. We investigate the relatedness of these studies to various properties of the Global Workspace Theory (GWT) in Appendix B.
@@
Studies on sparse Transformers have explored consolidating latent memory to extract localized, contextual representations from inputs [23, 24, 25, 26, 13, 27]. For example, Perceiver [26, 27] utilized iterative cross-attention with a latent array and transformation to capture dependencies across input tokens. Set Transformer [23] and Linear Unified Nested Attention [25] also employed iterative cross-attention but without the latent transformation. Other methods, like Blockwise Self-Attention, relied on a strong assumption of predefined modularization for attention specialization [28]. In contrast, our method learns such modularization and attention specialization through end-to-end training. Unlike the vast majority of methods that employ latent memory, our proposed method leverages both explicit memory and associative memory for enhanced contextual learning capability.
In a similar approach to constructing sparse Transformers based on GWT, the Coordination method [13] introduced a bottleneck into cross-attention mechanisms to encourage neural module specialization. The results indicated that the competition via the bottleneck could contribute to a small set of specialized priors in latent memory to facilitate relational learning. However, the number of priors was limited to fewer than 10, all with the same dimension as the input representation. By contrary, our method employs low-rank explicit memory to learn a significantly richer set of priors (up to 128 priors learned from a pool of 32.8k tokens). Moreover, the Coordination method relies on iterative cross-attentions, while our work focuses on novel associative memory-enabled sparse attention.
This work is also related to modular neural networks and mixture of experts [29, 30] in terms of competition in the shared workspace. Separating the information processing within Transformers into distinct components, depending on the input, has shown more flexibility in data processing.
@@
External memory such as tape storage and associative memory, has been employed in deep neural networks [31, 32, 33], with recent studies exploring the potential use of Hopfield networks and their modern variants [21, 34, 22]. In this work, we incorporate Hopfield networks as an integral element in the sparse attention mechanism of the proposed Associative Transformer. This approach is fundamentally different from the previous studies such as Energy Transformer [35] that mainly focused on utilizing Hopfield networks independently of the attention mechanism.
@@ 3 Associative Transformer
Figure 1: The scheme of the Associative Transformer (AiT). (a) In a global workspace layer, the input $\mathbb{R}^{B\times N\times E}$ is squashed into vectors $\mathbb{R}^{BN\times E}$. The squashed representations are projected to a latent space of dimension $D<<E$ and are sparsely selected to update the explicit memory via a fixed bottleneck $k<<BN$. The Hopfield network utilizes the memory to reconstruct the input tokens, where a learnable linear transformation (LT) scales the memory contents back to the input dimension $E$. (b) The Associative Transformer block consists of self attention, feed-forward layers, and the global workspace layer. Compared to Vision Transformer (ViT), leveraging the global workspace layer enhances the layer efficiency. A shallower 6-layer AiT is shown to outperform a 12-layer ViT (see Table 1).
This section discusses the key technical underpinnings of the Associative Transformer (AiT), focusing on the newly devised Global Workspace Layer (GWL). The GWL integrates three main elements: low-rank explicit memory, bottleneck attention, and associative memory (Figure 1). It facilitates efficient writing and reading in explicit memory by promoting competition among input tokens through a specialized top-$k$ bottleneck attention mechanism that limits cross-attention capacity. Two techniques are employed to maintain balanced information writing: Exponential Weighted Moving Average (EWMA), which stabilizes memory updates by combining old and new information, and a novel bottleneck attention balance loss, which encourages diverse token selection and counteracts the loss of specificity across layers. After writing, the explicit memory contents are broadcast back to the input tokens for retrieval within a Hopfield network-based associative memory, reconstructing the tokens into more globally meaningful representations that serve as the output of the GWL. In practice, the GWL can be integrated with conventional attention layers to enhance the efficiency of Transformer models, boosting their ability to learn abstract spatial relations across samples and time steps.
@@ 3.1 Vision Transformers for classification tasks
Vision Transformers (ViT) tackle image classification tasks by processing sequences of image patches. The pre-processing layer partitions an image into non-overlapping patches. Let $x\in\mathbb{R}^{H\times W\times C}$ be an input from the dataset $X$, where $(H,W)$ is the resolution of the image and $C$ is the number of channels. $x$ is separated into a sequence of patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(P,P)$ is the resolution of each image patch and $N=\frac{HW}{P^{2}}$ is the number of patches. These patches are mapped to embeddings $v_{p}\in\mathbb{R}^{N\times E}$ with the linear projection. ViT leverages self-attention to map a query and a set of key-value pairs to an output. The patch embeddings are used to obtain the query, key, and value based on linear transformations $W^{Q}\in\mathbb{R}^{E\times D},\,W^{K}\in\mathbb{R}^{E\times D},\text{ and, }W^{V}\in\mathbb{R}^{E\times D}$. The output of an individual attention head is a weighted sum of the values, $\mbox{softmax}(\frac{W^{Q}_{i}v(W^{K}_{i}v)^{T}}{\sqrt{D}})\,W^{V}_{i}v$.
@@ 3.2 Low-rank priors in explicit memory
To induce memory in large-scale models like Transformers, we focus on learning more efficient explicit memory, which is later consolidated with Hopfield networks. This allows us to decouple the memory storage from the Hopfield network learning, such that the memory could be gradually updated with old and new memory using methods such as exponentially weighted moving average (EWMA). The explicit memory is only trainable during model training and frozen during evaluation. The proposed Global Workspace Layer enables efficient writing and reading in the explicit memory to learn a set of low-rank priors through a tailored bottleneck attention mechanism.
In particular, we employ a squash layer to concatenate patch representations within the entire batch $V\in\mathbb{R}^{B\times N\times E}$ into vectors $\Xi\in\mathbb{R}^{BN\times E}$, fostering association among tokens across samples. In practice, the number of tokens $B\times N$ varies depending on the batch size $B$. The attention bottleneck with a fixed capacity $k$ restricts the number of attended tokens to a constant value. The output of the squash layer is then attended to by the multi-head cross-attention using the contents of the explicit memory as queries. The explicit memory is comprised of $M$ learnable low-rank priors $\gamma=\mathbb{R}^{M\times D}$ where $D$ is the dimension of the prior, $D<<E$. Priors are general assumptions about tokens to guide the attention for the token selection. The priors are utilized as various queries to compute cross-attentions that extract various sets of tokens and update the explicit memory. We found that a rank of 32 and initialization of the priors with a Gaussian distribution works well in most cases.
We devise a tailored cross-attention mechanism to update the explicit memory using the squash layer’s output $\Xi\in\mathbb{R}^{BN\times E}$. Notably, in the cross-attention, the query is a function of the current memory content $\gamma=\{\gamma_{i}\}_{i=1}^{M}$. The key and value are functions of the squash layer’s output $\Xi$. The attention scores $A_{i}$ for head $i$ are computed as follows $\text{A}_{i}(\gamma,\Xi)=\text{softmax}(\frac{\gamma(\Xi W^{K}_{i})^{T}}{\sqrt{D}})$. This is the case of soft attention without constraints on the bottleneck capacity.
@@ 3.3 Bottleneck attention with a limited capacity
We discuss the case of limiting the cross-attention capacity via top-$k$ bottleneck attention. The bottleneck attention constrains the number of tokens that can be attended to, fostering competition among tokens and ensuring the selection of the most relevant tokens. Notably, we select the tokens with the top-$k$ highest attention scores from $A_{i}$, ${h}_{i}=\text{top-}k(\text{A}_{i})\Xi W^{V}$. Additionally, to ensure a stable update of the explicit memory across different time steps, we employ layer normalization and the exponentially weighted moving average (EWMA) strategies as follows:
$\hat{\gamma}^{t+1}=(1-\alpha)\cdot\gamma^{t}+\alpha\cdot\text{LN}(\text{Concat}({h}_{1},\dots,{h}_{S})W^{O}),$ (2)
$\gamma^{t+1}=\frac{{\hat{\gamma}^{t+1}}}{{\sqrt{\sum_{{j=1}}^{M}(\hat{\gamma}^{t+1}_{j})^{2}}}},$
where $t$ is the batch time step, LN is the layer normalization, $S$ is the number of attention heads, and $\alpha$ is a smoothing factor determining the decay rate of older observations, which is a small value such as 0.1. EWMA ensures stable memory updates by accumulating old and new memories. We found that different batch sizes $B$ have little influence on the test performance. This can be attributed to the small decay rate that results in the stable memory update.
@@
The sparsity induced by the bottleneck attention results in the emergence of specialized priors by attending to various tokens. However, cascading multiple bottleneck attentions across Transformer layers could lead to difficulty in efficiently forming specialized priors. As information flows through multiple bottleneck attention layers, the representations become diluted, necessitating a mechanism to counteract this inherent loss of input specificity for the effective selection of meaningful tokens across layers. To overcome this challenge, we propose the bottleneck attention balance loss to encourage a more diverse token selection with the bottleneck attention. The bottleneck attention balance loss $\ell_{\text{bottleneck}}$ comprises two components, i.e., the cumulative attention loss $\ell_{\text{{importance}}}$, which computes the total assigned attention scores for each token, and the selected instance loss $\ell_{\text{{loads}}}$, which provides an estimate of the frequency of each token being selected. Then, the loss is computed for each token position $o\in\{1,2,\dots,B\times N\}$ in the squash layer output. We devise the bottleneck attention balance loss for attention head $i$ by measuring the normalized variances of the cumulative attention loss and the selected instance loss across token positions as follows:
$\ell_{\text{importance}_{i,o}}=\sum_{j=1}^{M}\text{A}_{i,j,o},\,\ell_{\text{loads}_{i,o}}=\sum_{j=1}^{M}(\text{A}_{i,j,o}>0),$ (3)
$\ell_{\text{bottleneck}_{i}}=\frac{{\text{Var}(\{\ell_{\text{importance}_{i,o}}}\}_{o=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{o=1}^{B\times N}\ell_{\text{importance}_{i,o}}})^{2}+\epsilon}+$ (4)
$\frac{{\text{Var}(\{\ell_{\text{loads}_{i,o}}}\}_{o=1}^{B\times N})}{{(\frac{1}{B\times N}\sum_{o=1}^{B\times N}\ell_{\text{loads}_{i,o}}})^{2}+\epsilon},$
where $\text{A}_{i,j,o}$ is head $i$’s attention score for the $j$th prior at token position $o$, Var($\cdot$) denotes the variance, and $\epsilon$ is a small value to avoid division by zero.
The losses for the different heads are summed up and added to the classification loss, $\ell=\ell_{\text{class}}+\sigma\cdot\sum_{i=1}^{S}\ell_{\text{bottleneck}_{i}},$ where $\sigma$ is a coefficient to adjust the ratio between the classification loss and the bottleneck attention balance loss.